a car starts form rest and accelerates uniformly over a time of 5.21 seconds for a distance of 110m. Determine the acceleration of the car


1
Expert's answer
2022-02-10T13:37:50-0500

Given:

"v_0=0"

"t=5.21\:\rm s"

"d=110\:\rm m"


The distance traveled by a car

"d=v_0t+at^2\/2=at^2\/2"

The acceleration

"a=2d/t^2=\rm 2*110\: m/(5.21\: s)^2=8.10\: m/s^2"